OpenStudy (anonymous):

What is the y−coordinate of the y−intercept of the line that passes through the points (−4, −4) and (4, 8)?

OpenStudy (anonymous):

Slope of the line is m = (8-(-4)) / ( 4-(-4)) = 12/8 = 1.5 For any straight line, y=mx+c, where m is the gradient and c is the y-intercept hence at point (4,8) 8 = 1.5*4 +c c = 8 - 6 =2 Hence the equation of the line passing through (-4,-4) and (4,8) will be y=1.5x+2 So the y-intercept will be 2

OpenStudy (anonymous):

Thank you soo much :)

OpenStudy (anonymous):

you're welcome =D